a sample of 9 g of methane (CH4) occupies a volume of 15.4 L at 37 C.
calculate the work done when the gas expands isothermally into a vacuuum space.
what are the equation we should use?
1
Expert's answer
2012-10-23T09:28:12-0400
w=-RT ln (V2/V1)=RT ln (P2/P1)

PV=nRT

n=9/16=0.5625
mol

P=nRT/V=0.5625*8.31*310 / 15.4 =94.095 kPa

w=-RT ln (V2/V1)=RT
ln (P2/P1)